### v2

# STC-ViT: Spatio Temporal Continuous Vision Transformer for Weather Forecasting

Abstract
Operational weather forecasting system relies on computationally expensive physics-based models. Recently, transformer based models have shown remarkable potential in weather forecasting achieving state-of-the-art results. However, transformers are discrete models which limit their ability to learn the continuous spatio-temporal features of the dynamical weather system. We address this issue with STC-ViT, a Spatio-Temporal Continuous Vision Transformer for weather forecasting. STC-ViT incorporates the continuous time Neural ODE layers with multi-head attention mechanism to learn the continuous weather evolution over time. The attention mechanism is encoded as a differentiable function in the transformer architecture to model the complex weather dynamics. We evaluate STC-ViT against a operational Numerical Weather Prediction (NWP) model and several deep learning based weather forecasting models. STC-ViT performs competitively with current data-driven methods in global forecasting while only being trained at lower resolution data and with less compute power.
1 Introduction
Operational weather forecasting systems rely on physics based Numerical Weather Prediction (NWP) models. While highly accurate, these system tend to be extremely slow with aggregated approximation errors Palmer et al. (2005) and requiring high computational resources usually hundreds of thousands of GPU nodes which is unfeasible Bauer et al. (2020) Andersson (2022). The recent transformers based Deep Learning (DL) architectures Nguyen et al. (2023a); Kurth et al. (2023); Nguyen et al. (2023b), have achieved impressive forecasting results outperforming the NWP models.
Figure 1: Conceptual representation of continuous dynamics of weather system entailing that the rate of change of air density in a given volume is balanced by the flux of air mass entering or leaving the volume.
Weather forecasting is based on the Continuity Equation $\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\textbf{u})=0$, which ensures the mass of air in the atmosphere remains constant over time, where $\rho$ is the air density, t is time, u is the velocity vector of the air, which has components u, v, and w representing the wind speeds in the x (east-west), y (north-south), and z (vertical) directions, respectively and $\nabla\cdot(\rho\textbf{u})$ represents the divergence of the mass flux.
However, transformers only process discrete sequences of input data and are fundamentally discrete in nature Fonseca et al. (2023) which limit their ability in modelling the continuous evolving dynamics of weather system.
In this work, we address the non-continuity problem in Vision Transformer (ViT) which arise in weather forecasting systems where data is present in discrete form. The distinctive and continuously evolving characteristics of weather data pose significant challenge in producing accurate forecasts. The correlation within weather information observed at different time steps is continuously evolving due to the continuous nature of data and processing it as discrete information greatly impacts the forecasting accuracy.
To tackle the temporal dynamics of a system, Neural Ordinary Differential Equation (NODE) Chen et al. (2018) was introduced. NODE is based on the concept of continuous depth modelling to learn smooth solutions from irregular time series data. However, while capturing the complex dependencies between observations, the iterative approach of Neural ODE’s can lead to escalating errors as the number of iterations increases Chen et al. (2018). Moreover, despite being computationally efficient and accurate, Neural ODEs are only limited to temporal data and fails to learn the continuous features of spatio-temporal dynamical systems such as weather forecasting.
To address these problems, we build upon the core ideas of Neural ODE Chen et al. (2018) and ViT Dosovitskiy et al. (2020). We propose STC-ViT which leverages the continuous learning paradigm to effectively learn the complex spatio-temporal changes even from weather data recorded at coarser resolution. The idea is to parameterize the attention mechanism by converting it into a differentiable function. Continuous temporal attention is calculated sample-wise and combined with the patch wise spatial attention to learn the spatio-temporal mapping of weather variables in the embedding space of the vision transformer. Furthermore, we add derivation as a pre-processing step to prepare the discrete data for continuous model and explore the role of normalization in continuous modelling.
Normalization is a well established concept in the field of deep learning especially in vision transformers where pre and post layer normalization optimize gradient propagation under different settings. The influence of different normalization techniques in neural ODEs was also explored by Gusak et al. however, the role of normalized derivatives is still poorly understood in complex continuous spatio-temporal data dependant forecasting systems. We empirically investigate the influence of normalization in forecasting different weather variables. We observed that normalizing the continuous evolution of hidden layers yields different results for each predicted weather variable. Empirical results show the ability of the proposed STC-Vit to model complex continuous spatio-temporal dynamic systems effectively and achieve state-of-the-art performance in global weather forecasting.
In this paper we focus on the following research question: Can we design a data-driven weather forecasting system that is computationally efficient, learns the continuous latent representation and achieves precision levels commensurate with Numerical Weather Prediction (NWP) model. In summary, our contributions are as follows:
Spatio-temporal continuous attention: We propose continuous attention to effectively learn and interpolate the spatio-temporal information for weather forecasting.
Data pre-processing for continuous systems: We introduce derivation as a pre-processing step to ensure better feature extraction for continuous spatio-temporal models.
Normalization in continuous modelling: We explore the role of normalization in STC-ViT and conduct extensive experiments to analyse our methodology.
2 Background
2.1 Attention in Vision Transformers.
The attention mechanism plays a crucial role in capturing long-term spatial dependencies. While the concept of attention in vision has been discussed in literature for decades, it gained popularity with the introduction of Vision transformers Dosovitskiy et al. (2020). The attention mechanism enables the model to focus on different parts of the input image simultaneously, capturing various features and their interactions at different positions. It allows the model to weigh the importance of different input features dynamically. In a transformer, attention is calculated using a set of queries $(Q)$, keys $(K)$, and values $(V)$, typically derived from the input features. The attention mechanism can be described mathematically as:
$Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
where Q, K, and V are matrices formed from all queries, keys, and values, respectively, $d_{k}$is the dimensionality of the keys and queries, and the division by $\sqrt{d_{k}}$ is a scaling factor to deal with gradient vanishing problems during training.
In practice, multi-head attention is used in which the attention mechanism is run several times in parallel allowing the model to jointly attend to information from different representation subspaces at different positions enhancing the model’s ability to represent complex dependencies.
2.2 Neural ODE
Neural ODEs are the continuous time models which learn the evolution of a system over time using Ordinary Differential Equations (ODE) Chen et al. (2018). The key idea behind Neural ODE is to model the derivative of the hidden state using a neural network. Consider a hidden state $h(t)$ at time $t$ In a traditional neural network, the transformation from one layer to the next could be considered as moving from time $t$ to $t+1$. In Neural ODEs, instead of discrete steps, the change in $h(t)$ over time is defined by an ordinary differential equation parameterized by a neural network:
$\frac{dh(t)}{dt}=f(h(t),t,\theta)$ (2)
where $\frac{dh(t)}{dt}$ is the derivative of the hidden state with respect to time, $f$ is a neural network with parameters $\theta$ and $t$ is the time variable, allowing the dynamics of $h(t)$ to change with time.
From ResNets to Neural ODEs To see the connection between ResNets and Neural ODEs, consider a ResNet with layers updating the hidden state as:
$h_{t+1}=h_{t}+f(h_{t},\theta_{t})$ (3)
In the limit, as the number of layers goes to infinity and the updates become infinitesimally small, this equation resembles the Euler method for numerical ODE solving, where:
$h(t+\triangle t)=h(t)+\triangle t\cdot f(h(t),t,\theta)$ (4)
Reducing $\triangle t$ to an infinitesimally small value transforms the discrete updates into a continuous model described by the ODE given earlier. To compute the output of a Neural ODE, integration is used as a backpropagation technique. This is done using numerical ODE solvers, such as Euler, Runge-Kutta, or more sophisticated adaptive methods, which can efficiently handle the potentially complex dynamics encoded by $f$.
3 Related Work
Integrated Forecasting System (IFS) ECMWF (2023) is the operational NWP based weather forecasting system which generates forecasts at a high resolution of 1km. IFS combines data assimilation and earth system model to generate accurate forecasts using high computing super computers. In contrast, data-driven methodologies have now outperformed NWP models with much less computational complexities.
WeatherBench Rasp et al. (2020) provides a benchmark platform to evaluate data-driven systems for effective development of weather forecasting models. Current state-of-the-art DL based models are mostly based on Graph Neural Networks (GNNs) and Transformer based models. Keisler Keisler (2022) implemented a message passing GNN based forecasting model which was further extended by GraphCast Lam et al. (2023) which implemented multi-mesh GNN to achieve state-of-the-art results. Pangu-Weather Bi et al. (2023) a 3D transformer based model was introduced which was designed to encode 3D earth pressure levels to reduce the forecasting errors. FourCastNet (FCN) Kurth et al. (2023) used a combination of Fourier Neural Operator (FNO) and ViT and was reported to be 80,000 times faster than NWP models. ClimaX Nguyen et al. (2023a) was introduced as a foundation model for weather and climate forecasting. The authors used Vision Transformer (ViT) as a backbone for spatio-temporal forecasting on number of tasks i.e. global forecasting, regional forecasting and seasonal forecasting each with a lead time of several hours, days and months.
Another data-driven model FengWu Chen et al. (2023a) showed remarkable performance on medium range weather forecasting. It uniquely approaches forecasting through a multi-modal and multi-task perspective, utilizing a deep learning architecture with model-specific encoder-decoders and a cross-modal fusion Transformer. Followed by FengWu, FuXi Chen et al. (2023b) emerged as high resolution forecasting system operating at $0.25^{\circ}$. FuXi integrates the data cube embedding with U-Transformer to produce 15-day forecasts. Recently, Stormer Nguyen et al. (2023b) based on transformer model emerged as the state-of-the-art weather forecasting model outperforming PanguWeather and GraphCast. While being highly accurate and showcasing remarkable scaling capabilities, these models are discrete space-time models and do not account for the continuous dynamics of weather system except for ClimODE Verma et al. (2023). ClimODE uses Neural ODE to incorporate a continuous-time process that models weather evolution and advection, enabling it to capture the dynamics of weather transport across the globe effectively, however it currently yields less precise forecast results compared to state-of-the-art models, offering significant potential for further enhancements.
4 Methodology
Problem Formulation. Consider a model $f$ receives weather data as input of the form $X^{N\times W\times H}$ at time $t$ and predicts the weather information at time $t+\triangle t$ as shown in equation 5 where $N$ is the number of weather variables such as temperature, wind speed, pressure, etc., $H\times W$ refers to the spatial resolution of the variable.
The objective of the model is to learn the continuous temporal dependencies between consecutive time steps while accounting for spatial correlations within the $H\times W$ grid. Since the weather changes continuously over time, it is essential to capture the continuous change within the provided fixed time step data. The main aim of STC-ViT is to to learn the continuous latent representation of the weather data using Spatio-Temporal Continuous Attention and Neural ODEs. The evolution of the weather system from $t$ to $t+\triangle t$ can be represented as:
$X(t+\triangle t)=f(X,t)$ (5)
$\frac{dX(t)}{dt}=f(X,t,\theta)$
$\int^{t+\triangle t}_{t}\frac{dX(t)}{dt}=\int^{t+\triangle t}_{t}f(X,t,\theta)$
$X(t+\triangle t)=[f(X,t,\theta)]^{t+\triangle t}_{t}$ $\theta=$ learnable model parameters
Derivation as a pre-processing step. Weather information is highly variational and complex at both temporal and spatial levels. Temporal derivatives of each weather variable are calculated to preserve weather dynamics and ensure better feature extraction from discretized data. We perform sample wise derivation at pixel level to capture the continuous changes in weather events over time as shown in equation 6.
$\frac{dV(x,y,t)}{dt}=\frac{V(x,y,t)-V(x,y,t-1)}{\Delta t}$ (6)
where $V(x,y,t)$ is pixel value of weather variable at time t and $V(x,y,t-1)$ is the pixel value of same variable at t-1. Figure 2 shows the graphical representation of the pre-processing step.
Figure 2: Temporal derivatives of an image time series capture the changes in pixel values over time. The process involves calculating the rate at which the intensity values of each pixel change from one time frame to the next.
Normalized Derivatives. Normalization of derivatives is one of the crucial steps to encode stability in the behavior of deep learning model. While the effects and results of different normalization techniques in neural ODEs has been explored before Gusak et al. (2020), it only looks at normalization as layers in the neural network architecture. We extend the idea by investigating the role of normalization when directly applied to derivatives. Normalizing derivatives involves scaling the derivative values so that they fall within a specific range or conform to a standard distribution. This normalization process can help improve the stability and performance of subsequent analysis. We applied two normalization most common techniques namely min-max and z-score normalization followed by the pre-derivation layer to showcase their benefits in continuous systems.
4.1 Spatio-Temporal Continuous Vision Transformer
Attention is one of the key components in the vision transformer architecture. It is based on the idea of determining the most crucial parts of image in the final prediction stage. Despite being successful in variety of tasks, ViT is still restrictive in its ability to learn informative embedding for a highly dynamical system like weather forecasting. We design multiple attention paths Temporal Continuous Attention (TCA) and Spatial Attention (SA) to model the the continuous evolving state of weather variables. The detailed architecture of STC-ViT is shown in figure 3.
Temporal Continuous Attention (TCA). First, we employ sample-wise attention to calculate the contextual embedding space for each weather variable evolving with time. This step ensures that in addition to spatial attention, the model also attends to the temporal evolution of the same variable across different samples within the batch. This temporal transformation is learned by assigning each variable with its own Query, Key, and Value across each input sample, similar to how it’s done within a single image. The attention mechanism calculates the attention scores between the variables of different samples (across the same variable positions).
This modification allows the model to capture relationships or dependencies between the same weather variables across different instances. It proved to be beneficial in weather forecasting scenario where the model is able to map the continuous evolving features of each weather variable. To ensure continuous learning, we introduce derivatives in the sample wise attention mechanism. Differential equations represent the dynamics of the physical system with respect to time and cater to the missing values of data. We combined the attention mechanism with the differential equation learning paradigm to model the change of the atmosphere for both the spatial and temporal features. Moreover, this approach removes the limitation of modelling complex physics equations in NWP models. Furthermore, instead of having predictions at particular time stamp only, STC-ViT learns the transitional changes from one time step to another which is important for capturing unprecedented changes in weather.
To compute TCA, we calculate derivatives of similarity across same variables of each data sample. Assume we have two input samples of size $(N\times W\times H)$ denoted as $X_{0}$ at time $t_{0}$ and $X_{1}$ at time $t_{1}$ where $N$ is the number of weather variables such as temperature, wind speed, pressure, etc., $H\times W$ refers to the spatial resolution of the variable. Each variable has its own Q, K, and V tensors across both samples. Given $Q_{t_{0},i}$ and $K_{t_{1},j}$ are the Query and Key representations for the $i_{th}$ variable at time $t_{0}$ and $j_{th}$ variable at time $t_{1}$. Continuous attention is computed as follows:
$\frac{d}{dt}(Q_{t_{0},i}).(K_{t_{1},j})=Q_{t_{0},i}\frac{d(K_{t_{1},j})}{dt}+K_{t_{1},j}\frac{d(Q_{t_{0},i})}{dt}$ (7)
$Q_{t_{0},i}\frac{d(K_{t_{1},j})}{dt}+K_{t_{1},j}\frac{d(Q_{t_{0},i})}{dt}=Y$ (8)
$Att(Q_{t_{0},i},K_{t_{1},j})=Softmax(\frac{Y}{\sqrt{d_{k}}})$ (9)
$TCA_{t_{0},t_{1}}=\sum\limits^{N}_{k=1}Att(Q_{t_{0},i},K_{t_{1},j})\cdot V_{t_{1},j}$ (10)
The resulting output represents the attention weighted sum of values for similar variables across input samples at time $t_{0}$ and $t_{1}$. This information captures relationships or dependencies between corresponding variables across the input samples at different time steps. This process calculates attention between similar variables across input samples at all time steps, allowing the model to capture relationships or interactions between variables across the entire sequence of input samples.
Spatial Attention (SA). For spatial attention, we simple calculate traditional attention scores along patches of the same image and obtain output as shown in equation 1 and call it SA.
Concatenation and Fusion. We concatenate the outputs of the TCA and SA along the feature dimension (as shown in equation 11) to provide a comprehensive representation, considering both temporal and spatial dependencies. This dual attention approach allows the Vision Transformer to effectively model complex spatio-temporal dependencies in the weather reanalysis data.
$STC=concat(TCA,SA)$ (11)
Neural ODEs as Add operation in Transformers. Neural ODEs model the transformation of data through a continuous dynamical system. Instead of discrete layers, the transformation is modeled by a differential equation. The idea is to treat the depth of the network as a continuous time variable t. As explained in section 2.2, Neural ODEs can be seen as a continuous form of ResNets.
In STC-ViT, instead of adding the output of an attention mechanism discretely as $h^{\prime}=LayerNorm(h+STC(h))$, we model the transformation using a Neural ODE. The attention output is treated as a continuous evolution of the input state $h$ transforming equation 2 as equation 12. Solving this ODE from $t=0$ to $t=1$ gives us the updated state $h(1)$, which incorporates the attention mechanism’s effect continuously over time.
$\frac{dh(t)}{dt}=STC(h(t),t,\theta)$ (12)
To compute the output of a Neural ODE, we solve the initial value problem:
[…]
This integral is computed using Runge-Kutta Runge (1895) numerical ODE solver. The solver adaptively determines the number of steps required, making the process efficient and accurate. This continuous approach allows for smoother transformations and improve model performance.
Figure 3: Whole prediction pipeline of STC-ViT as the model receives spatio-temporal weather information and passed through pre-derivation and normalization steps and input to transformer encoder where TCA and SA learns the continuous weather features and outputs the prediction
5 Experiments and Results
Dataset. We trained STC-ViT on ERA5 dataset at resolution of $5.625^{\circ}$ (32 x 64 grid points) and $1.40625^{\circ}$ (128 x 256 grid points) provided by WeatherBench (Rasp et al., 2020). Details of the dataset are provided in Appendix A.
Training Details. We consider weather forecasting as a continuous spatio temporal forecasting problem i.e a tensor of shape $N\times H\times W$ at time $t$ is fed to the pre-processing layer of the model where it passes through pre-derivation to STC-ViT and outputs a tensor of $N^{\prime}\times H^{\prime}\times W^{\prime}$ at future time step $t+\triangle t$. Complete training details of the model are given in the in Appendix B.
Normalization Experimental analysis show that z-score normalization produces superior results when applied to atmospheric variables such as geopotential at 500hpa pressure (z500) and temperature at 850hpa pressure (T850) and min-max produces better results when applied to surface variables such as 2-meter temperature (T2m) and mean sea-level pressure (MSLP). We compare the results of both methodologies at lead times of 1,3,5,7 and 10 days. The RMSE scores and comparisons are shown in figure 4.
Figure 4: The results comparison of normalized derivatives show that z-score normalization performs better for atmospheric variables while min-max produce better results for static variables
Ablation Studies We performed extensive empirical analysis to understand the role of different components in our model’s performance. The results of ablation studies in figure 5 show that while Neural ODE is a powerful architecture, it outperforms vanilla architectures when embedded in attention mechanism. Details of ablation studies are given in Appendix B.6.
Figure 5: Ablation studies highlight the results for different components of STC-ViT individually and show that the continuous attention component alone performs better than all other even NODE. Thus combined with NODE outperforms state-of-the-art results.
Results and Discussion We trained STC-ViT on ERA5 dataset at resolutions of $5.625^{\circ}$ for lead times ranging from 6 to 36 hours and $1.40625^{\circ}$ for lead times ranging from 1 to 10 days and call them $STC_{5.6}$ and $STC_{1.4}$ respectively. We compare $STC_{5.6}$ forecast results with ClimaX, FourCastNet and ClimODE on ERA5 dataset at $5.625^{\circ}$ resolution shown in Table 1. $STC_{5.6}$ outperforms ClimaX, FourCastNet and ClimODE at all lead times which shows that replacing regular attention with continuous attention in ViT architecture derives improved feature extraction by mapping the changes occurring between successive time steps.
Additionally, we compare the performance of $STC_{1.4}$ with the results of GraphCast and PanguWeather which are trained at a higher resolution of $0.25^{\circ}$ shown in Figure 7. $STC_{1.4}$ outperforms PanguWeather for medium-range forecasting starting at 3-day lead time and is on par with GraphCast. The superior performance of GraphCast and STC-ViT also opens a venue for future research in which image data could be modeled as graph and then processed through continuous attention mechanism. We also compare our results with both versions of IFS Andersson (2022), IFS-HRES which is state-of-the-art forecasting model at high resolution of $0.1^{\circ}$ and IFS-ENS, ensemble version trained at $0.2^{\circ}$.
Figure 6: RMSE comparison of STC-ViT trained at $1.40625^{\circ}$ with GraphCast and PanguWeather trained at $0.25^{\circ}$, IFS-ENS at $0.2^{\circ}$ and IFS-HRES at $0.1^{\circ}$ resolution data for lead times ranging from 1 to 10 days
Figure 7: RMSE results of $STC_{5.6}$ compared against ClimODE (ODE based), ClimaX (non-pretrained) and FourCastNet (FCN) trained on ERA5 at $5.625^{\circ}$ resolution
Table 1: RMSE and ACC results of $STC_{5.6}$ compared against ClimODE (ODE based), ClimaX (non-pretrained) and FourCastNet (FCN) trained on ERA5 at $5.625^{\circ}$ resolution
RMSE (Lower is better) ACC (Higher is better) Variable Lead Time (hrs.) STC-ViT ClimODE ClimaX FCN STC-ViT ClimODE ClimaX FCN z500 $(m^{2}\backslash s^{2})$ 6 59.61 102.9 247.5 149.4 1.0 0.99 0.97 0.99 12 76.20 134.8 265.3 217.8 1.0 0.99 0.96 0.99 18 86.10 162.7 319.8 275.0 1.0 0.98 0.95 0.99 24 92.09 193.4 364.9 333.0 1.0 0.98 0.93 0.99 36 111.31 259.6 455.0 449.0 0.99 0.96 0.89 0.99 T2m (K) 6 0.89 1.21 2.02 1.28 1.00 0.97 0.92 0.99 12 0.92 1.45 2.26 1.48 0.99 0.96 0.90 0.99 18 0.96 1.43 2.45 1.61 0.99 0.96 0.88 0.99 24 1.04 1.40 2.37 1.68 0.98 0.96 0.89 0.99 36 1.09 1.70 2.87 1.90 0.98 0.94 0.83 0.99 T850 (K) 6 0.84 1.16 1.64 1.18 1.00 0.97 0.94 0.99 12 0.88 1.32 1.77 1.47 1.00 0.96 0.93 0.99 18 0.92 1.47 1.93 1.65 0.99 0.96 0.92 0.99 24 1.11 1.55 2.17 1.83 0.99 0.95 0.90 0.99 36 1.19 1.75 2.49 2.21 0.98 0.94 0.86 0.99 U10 $(m\backslash s)$ 6 0.98 1.41 1.58 1.47 1.00 0.91 0.92 0.95 12 1.06 1.81 1.96 1.89 0.99 0.89 0.88 0.93 18 1.17 1.97 2.24 2.05 0.99 0.88 0.84 0.91 24 1.29 2.01 2.49 2.33 0.98 0.86 0.80 0.89 36 1.34 2.25 2.98 2.87 0.98 0.83 0.69 0.85
6 Conclusion and Future Work
In this paper, we present STC-ViT, a novel technique designed to capture continuous dynamics of weather data. STC-ViT achieves competitive results in weather forecasting which shows that vision transformers can model continuous nature of spatio-temporal dynamic systems with carefully designed attention mechanism.
While STC-ViT outperforms various state-of-the-art weather forecasting models, it is important to address its limitations in weather forecasting systems. STC-ViT is a deterministic model which does not account for model uncertainty which can produce unrealistic results. Extending STC-ViT to a probabilistic model can be addressed in future works. Also, scaling STC-ViT to better accommodate the multi-modal high resolution training and evaluation presents an opportunity as future research. Finally, addressing the black-box problem of STC-ViT can shed light on model learning insights which is equally important for climate science community.
Societal Impact
Our research focuses on modelling the continuous dynamics of weather forecasting system through the integration of deep learning (DL) techniques. The study shows that leveraging data-driven approaches can achieve significantly improved forecast results with compute efficient resources. The environmental benefits of compute-efficient forecasting systems are significant. Lowering the carbon footprint of computational processes contributes to global efforts in combating climate change. By integrating ML to improve accuracy while optimizing computational efficiency, we can create a sustainable and inclusive approach to weather forecasting that serves the global community more effectively.
Acknowledgements
This work was supported by SmartSAT Cooperative Research Center (project number P3-31s). We would also like to acknowledge NCI (National Computational Infrastructure) for providing us with GPU resources which enabled us to perform this research.
Appendix A Data
We used WeatherBench Rasp et al. (2020) to download the ERA5 Hersbach et al. (2018, 2020) dataset provided by the European Center for Medium-Range Weather Forecasting (ECMWF) for training our model. ERA5 is a weather reanalysis dataset of hourly intervals that provides data over several parameters such as land, ocean and atmosphere. The real-time data with the highest resolution of $0.25^{\circ}\times 0.25^{\circ}$ is being collected since 1940 till date. ERA5 is considered to be the most physically accurate dataset as it is based on data assimilation methodology in which new forecasts are continuously combined with previous forecasts in such a way that resultant is consistent with laws of physics. WeatherBench provides regridded data at resolutions of $5.625^{\circ}$ , $2.8125^{\circ}$ and $1.40625^{\circ}$. Table 2 provides a detail of weather variables used for training STC-ViT. For our experiments we split the data into training (1979-2017), validation (2018) and test (2019-2021) sets.
Table 2: ERA5 dataset variables and their atmospheric pressure values
Variable name (Type) Abbreviation Levels Geopotential (Atmospheric) Z 50 250 500 600 700 850 925 U Wind Component (Atmospheric) U 50 250 500 600 700 850 925 V Wind Component (Atmospheric) V 50 250 500 600 700 850 925 Temperature (Atmospheric) T 50 250 500 600 700 850 925 Specific Humidity (Atmospheric) Q 50 250 500 600 700 850 925 Relative Humidity (Atmospheric) R 50 250 500 600 700 850 925 Land-sea mask (static) LSM Orography (static) 2 meter Temperature (static) T2m Mean Surface level Pressure (static) MSLP 10 meter U Wind Component (static) U10 10 meter V Wind Component (static) V10
Appendix B Experiment Details
B.1 Evaluation Metrics.
We used Root Mean Square Error (RMSE) and Anomaly Correlation Coefficient to evaluate our model predictions. The formula used for RMSE is:
$RMSE=\frac{1}{N}\sum\limits^{N}_{k=1}\sqrt{\frac{1}{H\times W}\sum\limits^{H}_{i=1}\sum\limits^{W}_{j=1}L(i)(\hat{X}_{k,i,j}-X_{k,i,j})^{2}}$ (14)
where $H\times W$ is the spatial resolution of the weather input and N is the number of total samples used for training or testing. L(i) is used to account for non-uniformity in grid cells.
$ACC=\frac{\sum_{k,i,j}\hat{X^{\prime}}_{k,i,j}X^{\prime}_{k,i,j}}{\sqrt{\sum_{k,i,j}L(i)\hat{X^{\prime}}^{2}_{k,i,j}\sum_{k,i,j}L(i)X^{\prime 2}_{k,i,j}}}$ (15)
Where $\hat{X^{\prime}}=\hat{X^{\prime}}-C$ and $X^{\prime}=X^{\prime}-C$ and C is the temporal mean of the entire test set $C=\frac{1}{N}\sum_{k}X$
B.2 Optimization
We use the AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of 0.001. We utilize a 10% linear warm-up, after which a Cosine learning rate scheduler is adopted which progressively lowers the learning rate to zero following the warm-up period. We train STC-ViT for 50 epochs which takes about 4-5 days in total.
B.3 Hyperparameters
Table 3: Hyperparameters of STC-ViT
Hyperparameters Meaning Value
N Number of variables 48
p Patch size 4
Heads Number of continuous attention heads 16
Depth Number of Transformer layers 8
Dimension Hidden dimensions 1024
Dropout Dropout rate 0.1
batch_size Batch Size 4
$\beta 1$ First Moment Decay Rate of AdamW optimizer 0.9
$\beta 2$ Second Moment Decay Rate of AdamW optimizer 0.999
ES Early stopping True
ES rate Early stopping tolerance 20
B.4 Normalization
In our experiments, we normalize all inputs during training and re-scale them to their original range before computing final predictions. We perform z-score normalization for every variable, at each atmospheric pressure level, we calculate the mean and standard deviation to standardize them to zero mean and unit variance.
B.5 Software and Hardware Requirements
We use PyTorch Paszke et al. (2019), Pytorch Lightning Falcon (2019), torchdiffeq Chen et al. (2018), and xarray Hoyer and Hamman (2017) to implement our model. We use 8 DGXA100 devices with 80GB VRAM for training STC-ViT. All architectural and training details are provided at https://anonymous.4open.science/r/STC-ViT-3276.
B.6 Ablation Studies
Continuous Attention only model. To understand the importance of continuous attention, we remove the neural ODE layer and provide the output of attention mechanism to the basic feed forward network. While continuous attention only model does not outperform the STC-ViT, it provides better prediction accuracy than the traditional attention model.
Continuous Attention vs Neural ODE layer. We perform another ablation study to compare how well Neural ODE capture continuity in weather information when simply added as a layer in the transformer architecture. Neural ODE Chen et al. (2018) is a continuous depth neural network designed to learn continuous information in the process as well. For this study, we replace the continuous attention with vanilla attention and add a Neural ODE layer in the feed forward block. The feed forward solution is approximated using Runge-Kutta (RK) numerical solver.
Vanilla Vision Transformer Vision Transformer has emerged as a powerful architecture which captures long-term dependencies better than any model. However, this ablation study shows the limitation of ViT in representing the dynamical systems such as weather forecasting. For this ablation study, we simply train a basic ViT architecture on ERA5 dataset. Compared with STC-ViT, ViT under performs for prediction accuracy showing the superiority of continuous models in weather forecasting systems.
Vanilla Neural ODE Network. Another ablation study is done using vanilla Neural ODE model. We replace the transformer with Neural ODE architecture as proposed in the original paper Chen et al. (2018). While Neural ODEs are computationally efficient, they only aid in interpolating the temporal irregularities and ignore the spatial continuity. This study proves that STC-ViT learns both spatio-temporal continuous features from the discrete data and is better at representing dynamical forecast systems.
Appendix C Results Summary
Figure 8: 1 day forecasts
Figure 9: 3 day forecasts
Figure 10: 5 day forecasts